OpenStudy (anonymous):

Translate this real-world problem into an equation: Jenna's dog Bruiser eats 5 times the number of ounces of food that her dog Princess eats. Bruiser also eats 4 milk-bones, for a total of nineteen ounces of food. How much food does Princess eat?

OpenStudy (anonymous):

5x+4=19....

OpenStudy (anonymous):

@Chocolateo2

OpenStudy (perl):

lets call the regular food that bruiser and princess eat , B and P then B= 5*P now the non regular food (milkbones) is 4 ounces. so we have 5p+ 4 = 19